OpenStudy (anonymous):

I cant find this answer it is a quadratic equation 4/x+8/x+4=2

OpenStudy (anonymous):

(4/x)+(8/x+4) = 2 (4(x+4) + 8x)/x(x+4) = 2 4x+16+8x / x^2 +4x =2 12x+16 = 2(x^2 +4x) 12x+16=2x^2+4x 0 = 2x^2+4x-12x-16 0 = 2x^2-8x-16

OpenStudy (anonymous):

what does 4/x mean?